OpenStudy (anonymous):

The sum of two consecutive odd integers is 236. What is the smaller integer?

OpenStudy (zarkon):

236/2-1

OpenStudy (anonymous):

x + x + 2 = 236 2x + 2 = 236 2x = 234 x = 117

OpenStudy (anonymous):

1st: 2x-1 2nd: 2x+1 4x = 236 x = 59 : 2x-1 = 117 : 2x+1 = 119

OpenStudy (anonymous):

then solve for x